### v3

@@ Abstract
Text-to-image diffusion models allow users control over the content of generated images. Still, text-to-image generation occasionally leads to generation failure requiring users to generate dozens of images under the same text prompt before they obtain a satisfying result. We formulate the lottery ticket hypothesis in denoising: randomly initialized Gaussian noise images contain special pixel blocks (winning tickets) that naturally tend to be denoised into specific content independently. The generation failure in standard text-to-image synthesis is caused by the gap between optimal and actual spatial distribution of winning tickets in initial noisy images. To this end, we implement semantic-driven initial image construction creating initial noise from known winning tickets for each concept mentioned in the prompt. We conduct a series of experiments that verify the properties of winning tickets and demonstrate their generalizability across images and prompts. Our results show that aggregating winning tickets into the initial noise image effectively induce the model to generate the specified object at the corresponding location.
@@ 1 Introduction
For better theoretical understanding of diffusion models. The study of winning tickets offers a pathway to deepen our understanding of the denoising process, potentially benefiting research on diffusion model training and sampling.
@@ 3.1 Cross-Attention Layer and Winning Tickets
Cross-Attention Layer The cross-attention maps can serve as a monitor displaying the similarity between each pixel block and each concept mentioned in the text prompt [24]. Stable diffusion augments the U-Net [33] with the cross-attention mechanism [36] to allow for flexible condition. For the text-to-image task, a CLIP is employed to project text $p$ to a representation $\tau_{\theta}(p)$, which is subsequently mapped to the intermediate layers of the U-Net through a cross-attention layer as follows,
@@
Winning Ticket Following [24], we partition each random noise image, sized $64\times 64$, into $4\times 4$ pixel blocks, resulting in a total of $256$ blocks per image. Each attention value on the map $M^{(i)}$ is determined by a $4\times 4$ pixel block on the noisy image $z_{t}$ and a word embedding $p^{(i)}$, signifying the resemblance between this block and the concept $p^{(i)}$. Utilizing this attention map, the cross-attention layer reinforces features of $p^{(i)}$ into areas already resembling it, facilitating the denoising of pixel blocks initially exhibiting high cross-attention values for concept $p^{(i)}$. Pixel blocks demonstrating high cross-attention values on the map corresponding to word $p^{(i)}$ are identified as the winning tickets for concept $p^{(i)}$.
@@ 3.2 Lottery Tickets Collection
In this section, we present a method to collect all pixel blocks (lottery tickets without knowing if they won) along with their category-wise similarity. To construct this database, a list of possible generation categories $\mathcal{L}$ needs to be specified in advance. Specifically, for the given list of possible categories $\mathcal{L}$, we enumerate all possible combinations of two different objects, $\{c_{1},c_{2}\}$, to construct a list of prompts $\mathcal{L}_{p}$. The prompts take the form of [a $c_{1}$ and a $c_{2}$].
In practical, we find that when the prompt mentions only one object (e.g., a photograph of a $c$), the pixel blocks with high attention values for the word $c$ are winning tickets for "foreground" rather than winning tickets for the object $c$, due to the absence of competition between different concepts. Therefore, in our experiments, we utilize contrastive prompts, i.e., prompts composed of a pair of objects (e.g., [a $c_{1}$ and a $c_{2}$]), and perform a single step of denoising through the U-Net in the model. We collect the similarity information between tickets and these two concepts according to the cross-attention maps, as shown in the upper part of Fig. 2. Since the attention values of blocks are obtained through soft-max over words, obtaining a high score for word $c_{1}$ implies that this pixel block is a winning ticket for word $c_{1}$ over word $c_{2}$. We believe that such a design enables the screening of winning tickets with significant semantic information.
@@
$\text{Average Score}_{i}=\frac{1}{\mathcal{N}_{c}-1}\sum_{j\neq i}\mathcal{S}_{i,(j)}$ (2)
@@ 3.3 Winning Tickets Selection
$\mathcal{B}_{i}=\{t~{}|~{}\text{Average Score}_{i}(b)>\mathcal{T}_{obj}\},$ (4)
where $\text{Average Score}_{i}(b)$ denotes the average score of the pixel block $b$ on the category $c_{i}$. We select pixel blocks with low average scores for all categories $c_{i}\in\mathcal{C}$ to obtain winning tickets for the background.
@@ 4.1 Semantic-Driven Initial Image Construction
Figure 3: The collection contains scores for all lottery tickets, contrasting each category against others ($c_{i}$ (vs. $c_{j}$)). These scores serve as critical metrics used in the selection.
@@ 4.2 Collection On the Fly
In this section, we manually create prompts containing multiple objects and manually specifying the locations of the individual objects to confirm the effectiveness of winning tickets in application. After we are given the prompt $p$, we use the set of objects contained in the prompt as the possible category list $\mathcal{L}$ and select the winning tickets from multiple random noise images. As shown in Fig. 4, and 5, selected winning tickets proficiently direct the generation of the three specified objects. Moreover, if descriptive words related to the background, such as sky, lake, etc., are included in the categories list during the collection creation, we can also find winning tickets for these background elements. In Fig. 6, the top and bottom halves utilize identical prompts, differing only in the designated background regions. This difference is prominently reflected in the denoising results, demonstrating the power of winning tickets.
@@ 4.3.3 Combination with related methods
Some training-free mask-based methods [3, 23] are proposed for the layout-to-image synthesis task. We compare our method to them, and also since these methods focus on providing additional guidance merely through adding masks on cross-attention maps during the denoising process, they are fully orthogonal to what is discussed in this paper. We try to use our method in conjunction with these methods to confirm that initialization using winning tickets improves the performance of these related methods. A recent work [24] proposes to manipulate initial noise inside only a single noise image but suffers from the limited number of winning tickets. Therefore, we also try to improve its performance by combining them with ours.
@@ 4.3.5 Hyper Parameters
Table 1: Constructing initial noise solely through the utilization of winning tickets achieves a control effectiveness comparable to existing layout-to-image methods, and incorporating winning tickets alongside them significantly enhances their performance.
@@ 4.4 Discussion
The successful generation and the ability to exert control through utilizing winning tickets are enlightening for three significant reasons as follows:
@@
Figure 7: Versatility of winning tickets over different prompts containing the same objects. These images are generated from the same constructed noise and different prompts. The tickets constructing each area are selected according to their scores on the simple prompt a dog and a car. The generation tendency of those tickets remains unchanged even with different descriptions, such as black or running dog.
@@ 4.5 Limitations and Future Works
The position control of the third image is successful, but the quality of this image is compromised and the background generation fails. As a small number of the initial noise constructed by winning tickets deviates extremely from the normal distribution, images generated from them carry certain unnatural features. Recalibrating the constructed initial image to adhere to a normal distribution may alleviate this issue. The last image contains two very small objects, and our method only succeeds in generating the person. Although our method significantly outperforms other methods on control over small objects, when the guidance box is extremely small, the generation is still challenging.

### v4

@@ Abstract
Text-to-image diffusion models allow users control over the content of generated images. Still, text-to-image generation occasionally leads to generation failure requiring users to generate dozens of images under the same text prompt before they obtain a satisfying result. We formulate the lottery ticket hypothesis in denoising: randomly initialized Gaussian noise images contain special pixel blocks (winning tickets) that naturally tend to be denoised into specific content independently. The generation failure in standard text-to-image synthesis is caused by the gap between optimal and actual spatial distribution of winning tickets in initial noisy images. To this end, we implement semantic-driven initial image construction creating initial noise from known winning tickets for each concept mentioned in the prompt. We conduct a series of experiments that verify the properties of winning tickets and demonstrate their generalizability across images and prompts. Our results show that aggregating winning tickets into the initial noise image effectively induce the model to generate the specified object at the corresponding location. Project Page: https://ut-mao.github.io/noise.github.io
@@ 1 Introduction
For better theoretical understanding of diffusion models. The study of winning tickets offers a pathway to deepen our understanding of the denoising process, potentially benefiting research on diffusion model training and sampling acceleration.
@@ 3.1 Cross-Attention Layer and Winning Tickets
Cross-Attention Layer The cross-attention maps can serve as a monitor displaying the similarity between each pixel block and each concept mentioned in the text prompt [24]. Stable diffusion augments the U-Net [33] with the cross-attention mechanism [36] to allow for flexible condition. For the text-to-image generation task, a CLIP is employed to project text $p$ to a representation $\tau_{\theta}(p)$, which is subsequently mapped to the intermediate layers of the U-Net through a cross-attention layer as follows,
@@
Winning Ticket In this study, we leverage attention maps from the most down-sampled cross-attention layer, which effectively preserves pixel block information [24]. We partition each random noise image, sized $64\times 64$, into $4\times 4$ pixel blocks, resulting in a total of $256$ blocks per image. Each attention value on the map $M^{(i)}$ is determined by a $4\times 4$ pixel block on the noisy image $z_{t}$ and a word embedding $p^{(i)}$, signifying the resemblance between this block and the concept $p^{(i)}$. Utilizing this attention map, the cross-attention layer reinforces features of $p^{(i)}$ into areas already resembling it, facilitating the denoising of pixel blocks initially exhibiting high cross-attention values for concept $p^{(i)}$. Pixel blocks demonstrating high cross-attention values on the map corresponding to word $p^{(i)}$ are identified as the winning tickets for concept $p^{(i)}$.
@@ 3.2 Lottery Tickets Collection
In this section, we present a method to collect all pixel blocks (lottery tickets without knowing if they won) along with their category-wise similarity. To construct this database, a list of possible generation categories $\mathcal{L}$ needs to be specified in advance. Specifically, for the given list of possible categories $\mathcal{L}$, we enumerate all possible combinations of two different objects, $\{c_{1},c_{2}\}$, to construct a list of prompts $\mathcal{L}_{p}$. The prompts take the form of [a $c_{1}$ and a $c_{2}$] (e.g. [a dog and a cat]).
In our practical experience, we find that when the prompt mentions only one object (e.g., a photograph of a $c$), the pixel blocks with high attention values for the word $c$ are winning tickets for "foreground" rather than winning tickets for the object $c$, due to the absence of competition between different concepts. Therefore, in our experiments, we utilize contrastive prompts, i.e., prompts composed of a pair of objects (e.g., [a $c_{1}$ and a $c_{2}$]), and perform a single step of denoising through the U-Net in the model. We collect the similarity information between tickets and these two concepts according to the cross-attention maps, as shown in the upper part of Fig. 2. Since the attention values of pixel blocks are obtained through soft-max over words, obtaining a high score for word $c_{1}$ necessarily implies that this pixel block is a winning ticket for word $c_{1}$ over word $c_{2}$. We believe that such a design enables the screening of winning tickets with significant semantic information.
@@
$\vspace{-5pt}\text{Average Score}_{i}=\frac{1}{\mathcal{N}_{c}-1}\sum_{j\neq i}\mathcal{S}_{i,(j)}$ (2)
@@ 3.3 Winning Tickets Selection
$\vspace{-5pt}\mathcal{B}_{i}=\{t~{}|~{}\text{Average Score}_{i}(b)>\mathcal{T}_{obj}\},$ (4)
where $\text{Average Score}_{i}(b)$ denotes the average score of the pixel block $b$ on the category $c_{i}$.
We select pixel blocks with low average scores for all categories $c_{i}\in\mathcal{C}$ to obtain winning tickets for the background.
@@ 4.1 Semantic-Driven Initial Image Construction
Figure 3: The collection contains scores for all lottery tickets, contrasting each category against others ($c_{i}$ (vs. $c_{j}$)). These scores serve as critical metrics used in the selection of the winning ticket.
@@ 4.2 Collection On the Fly
In this section, we manually create prompts containing multiple objects and manually specifying the locations of the individual objects to confirm the effectiveness of winning tickets in application. After we are given the prompt $p$, we use the set of objects contained in the prompt as the possible category list $\mathcal{L}$ and select the winning tickets from multiple random noise images. As shown in Fig. 4, and 5, selected winning tickets proficiently direct the generation of the three specified objects. Moreover, if descriptive words related to the background, such as sky, lake, etc., are included in the categories list during the collection creation, we can also find winning tickets for these background elements. In Fig. 6, the top and bottom halves utilize identical prompts, differing only in the designated background regions. This difference is prominently reflected in the denoising results, demonstrating the power of winning tickets. More visual samples are provided in the supplementary.
@@ 4.3.3 Combination with related methods
Some training-free mask-based methods [3, 23] are proposed for the layout-to-image synthesis task. We compare our method to them, and also since these methods focus on providing additional guidance merely through adding masks on cross-attention maps during the denoising process, they are fully orthogonal to what is discussed in this paper. We, therefore, try to use our method in conjunction with these methods to confirm that initialization using winning tickets improves the performance of these related methods. A recent work [24] proposes to manipulate initial noise inside only a single random noise image but suffers from the limited number of winning tickets. Therefore, we also try to improve its performance by combining them with ours.
@@ 4.3.5 Hyper Parameters
Table 1: Constructing initial noise solely through the utilization of winning tickets achieves a control effectiveness comparable to existing layout-to-image methods, and incorporating winning tickets alongside them significantly enhances their performance further.
@@ 4.4 Discussion
The successful generation of images and the ability to exert effective control through utilizing winning tickets are profoundly enlightening for three significant reasons as follows:
@@
Figure 7: Versatility of winning tickets over different prompts containing the same objects. These four images are generated from the same constructed initial image and different prompts. The pixel blocks used for constructing each area are selected according to their scores on the simple prompt a dog and a car. The generation tendency of those pixel blocks remains unchanged even with different descriptions, such as black dog or running dog.
@@ 4.5 Limitations and Future Works
The position control of the third image is successful, but the quality of this image is compromised and the background generation fails. As a small number of the initial noise constructed by winning tickets deviates extremely from the normal distribution, images generated from them carry certain unnatural features. Recalibrating the constructed initial image to adhere to a normal distribution may alleviate this issue.
The last image contains two very small objects, and our method only succeeds in generating the person. Although our method significantly outperforms other methods on control over small objects, when the guidance box is extremely small, the generation is still challenging.